# Optimal Spatiotemporal Resolutions

### v2

@@ Abstract
In general, the comprehension of any type of complex system depends on the resolution used to look at the phenomena occurring within it. But identifying a priori, for example, the best time frequencies/scales to study a certain system over-time, or the spatial distances at which looking for correlations, symmetries, and fluctuations, is most often non-trivial. Here e describe an unsupervised approach that, starting solely from the data of a system, allows learning the characteristic length-scales of the key events/processes dominating it and the optimal spatiotemporal resolutions to characterize them. We test this approach on time-series data obtained from simulation or experimental trajectories of various example many-body complex systems ranging from the atomic- to the macroscopic-scale and having diverse internal dynamic complexities. Our method automatically performs the analysis of the systems’ data analyzing correlations at all relevant inter-particle distances and at all possible inter-frame intervals in which their time-series can be subdivided: namely, at all space-and-time resolutions. The optimal spatiotemporal resolution for studying a certain system thus steps-out as that maximizing information extraction-and-classification from the system’s data, which we prove being related to the characteristic spatiotemporal length-scales of the local/collective physical events dominating it. This approach is broadly applicable and can be used to optimize the study of different types of data (static distributions, time-series, or signals). The concept of “optimal resolution” has general character and provides a robust basis to characterize any type of system based on its data, as well as to guide data analysis in general.
@@ Introduction
Understanding the physics of complex many-body systems with rich internal dynamics is typically a challenging task [ten_wolde_enhancement_1997, baletto_structural_2019, gasparotto_identifying_2020, bochicchio_how_2019, de_marco_controlling_2021, cioni_innate_2023, crippa_detecting_2023]. This holds true across various scales [cho_dynamics_2021], from complex phenomena at the molecular level, such as nucleation events, phase transitions and ensemble dynamic adaptations in complex molecular systems [ten_wolde_enhancement_1997], to non-trivial collective behaviors in complex scale systems such as fish schools [porfiri_inferring_2018, butail_model-free_2016], bird flocks [cavagna_scale-free_2010, attanasi_information_2014, nagy_hierarchical_2010] and socioeconomic or stock market systems [liu_quantifying_2021, mantegna_introduction_1999, duan_network_2022]. Typically, a first challenge in studying such complex systems is to extract relevant information about the natural dynamical behavior of their constituent units, a fundamental step in reconstructing the complex internal physics that characterize them [martino2024data].
@@
The study of a complex system typically begins with the acquisition of trajectories – such as positions and velocities – of the individual units composing the system, obtained either through experiments or simulations. From these trajectories, an efficient method is needed to extract physically relevant information. A common approach involves using descriptors, which convert raw trajectories into time-series that can represent functions of positions (e.g., mutual arrangements) and/or velocities of the constituent units over time [kathirgamanathan_feature_2020, schmidt_human-based_2023]. The choice of descriptor is critical, as different descriptors enable the extraction of different types of physical information about the system’s behavior [caruso_timesoap_2023, crippa_molecular_2022, martino2024data]. Among the variety of available descriptors, a distinction can be made between physics-inspired descriptors and “abstract” or general ones. Physics-inspired descriptors are tailored to study specific systems, leveraging prior knowledge of their physics. While they can be highly effective for investigating particular types of systems, they are typically less transferable to others. Furthermore, these descriptors may suffer from inherent biases, potentially overlooking important and unforeseen information. To address these limitations, more general “abstract” descriptors – such as correlation functions or relative distances, to name a few commonly used examples – can be employed.
@@
More advanced and complex abstract descriptors, such as the Smooth Overlap of Atomic Positions (SOAP) [bartok_representing_2013] and Atomic Cluster Expansions (ACE) [drautz_atomic_2019] have also proven to be highly useful. Unlike system-specific descriptors, these general approaches are not tailored ad hoc to any particular system, making them applicable to virtually any type of system. Their unbiased nature, stemming from the lack of reliance on prior knowledge of the system’s physics, allows them to be broadly applicable and transferable for studying a wide range of complex molecular systems. Pattern recognition approaches for analyzing datasets, such as those from ACE or SOAP, have successfully provided valuable insights into the physics of systems across scales, from molecular to macroscopic, including aqueous, metallic, and other complex systems [fitzner2019ice, Capelli2022Ephemeral, lionello2022supramolecular, cioni_innate_2023, donkor_machine-learning_2023, cioni_innate_2023, crippa_machine_2023, cioni2024sampling, donkor_beyond_2024, martino2024data, perrone2024unsupervised, cheng_ab_2019, rapetti_machine_2023], and large-scale phenomena such as wave propagation [becchi_layer-by-layer_2024, caruso2024classification]. Data mining approaches often rely on identifying high-density peaks (dominant microscopic structural motifs) in datasets, treating the data as an ensemble where time correlations are ignored. This can potentially lead to a loss of information [martino2024data, lionello2024relevant].
@@
Conversely, it has been demonstrated how studying data, acquired over time-series (thus accounting for their time-correlation) allows decouple temporal from spatial information, which may provide more rich information than typical pattern-recognition approaches, including details on rare events, local or collective events that originate and amplify dynamically over time, etc [caruso_timesoap_2023, crippa_detecting_2023, butler_change_2024, crippa_machine_2023, becchi_layer-by-layer_2024, cioni2024sampling, perrone2024unsupervised, martino2024data, caruso2024classification, lionello2024relevant, husic_markov_2018]. For example, descriptors such as timeSOAP [caruso_timesoap_2023] or LENS [crippa_detecting_2023] track how much and in what way the environment surrounding each unit in the system changes over time. These approaches shift the focus to identifying relevant fluctuations in noisy time-series. An efficient method for this is Onion Clustering, an unsupervised and parameter-free algorithm that systematically detects statistically significant fluctuations in noisy time-series of any type. It classifies these fluctuations based on their similarity, determining how many distinct microscopic domains (clusters) can be identified in a time-series as a function of the length of the signal intervals used in the analysis. Recent results from Onion clustering have revealed non-trivial dependencies on the temporal resolution used in the analysis. These findings show that, for a given system studied with a specific descriptor, there exists an optimal temporal resolution that maximizes the information extracted from its time-series data [becchi_layer-by-layer_2024, martino2024data, lionello2024relevant]. The considerations for temporal resolution also apply to spatial resolution. More broadly, this raises an important question: what is the optimal resolution – both spatial and temporal – for capturing the physical behavior of a system? For instance, it is known that adjusting the cutoff for radial-based local descriptors (e.g., capturing different ranges of short- or long-range effects) can significantly impact analysis results. Similar to the findings from Onion Clustering regarding temporal resolution, is there an optimal spatial resolution for extracting meaningful physical information from trajectory data? Moreover, is there a reliable method to determine this optimal resolution for a given system?
@@
In this work, we address this fundamental problem by studying various prototypical systems, ranging from atomic and molecular scales to microscopic and macroscopic scales, which exhibit different complex internal dynamics. We systematically analyze the information contained in the time-series data of their constituent units as a function of the spatial and temporal resolution used in the analyses. We demonstrate that different types of systems have distinct optimal spatiotemporal resolutions maximizing the extraction of relevant information from their trajectories, depending on the local or collective nature of the physical events/processes that occur within them. The characteristic spatiotemporal length scales of the dominant physical events dominating the system are in this way revealed in a purely data-driven way, and without prior assumptions on the best resolutions to capture and characterize them.
@@
While identifying the optimal time and space resolutions to study a system about which little is known a priori is challenging, this work demonstrates how such information is intrinsically present in the system’s data and it provides an efficient way to extract it. From a fundamental point of view, we provide clear demonstrations of the impact that such a data-driven approach can have in revealing the physics of complex systems (or phenomena) without relying on prior knowledge or preconceived assumptions. At the same time, since every set of data has its own optimal spatiotemporal resolutions that allow for maximizing information extraction out of it, we obtain a general concept and an useful approach to guide and assist the analysis of virtually any type of data.
@@ Results
Figure 1: Extracting information from, e.g., LENS time-series of ice/water dynamic coexistence simulation trajectories. A Scheme of example local dynamical events captured by the LENS descriptor (permutation, addition, or subtraction of neighbors). B LENS signals for all the water molecules (their oxygen atoms) in the system as a function of simulation time, and their cumulative distribution (KDE: on the right, in blue). C Same LENS time-series data with the background colored based on the three main micro-clusters detected by Onion Clustering (using a time-resolution of $\Delta t=1.1$ ns). The gaussian LENS environments are shown in the KDE (right) as solid gaussian curves, the inter-cluster thresholds are indicated as dotted horizontal lines. D Output Onion plot. The curve in blue (primary $y$-axis) shows the number of clusters classifiable by Onion Clustering as a function of the time resolution $\Delta t$. The curve in orange (secondary $y$-axis) shows the fraction of unclassifiable data points (stored into a cluster named ENV0) as a function of $\Delta t$ (i.e., fraction of dynamical events occurring faster than the resolution of the analysis). The vertical red dashed line indicates the time-resolution $\Delta t=1.1$ ns used for the other panels (clusters and thresholds in panel C). E MD snapshot of the ice/water coexistence simulation showing the TIP4P/ICE molecules (left) and coloring them based on the clusters classified at the example time-resolution, which correspond to the bulk of ice (in gray), to the bulk of the liquid phase (in blue), and to the ice/water interface (in red). In white are the unclassifiable points: these are domains in the liquid which freeze and re-melt faster than $\Delta t=1.1$ ns [becchi_layer-by-layer_2024, Capelli2022Ephemeral, crippa_detecting_2023, caruso_timesoap_2023]. F Schematic representation of the solvation shells accounting for the first, second, third, etc., neighbor particles around a unit i in a generic system: using different cutoff radii means capturing different types of events (e.g., local vs. non-local) whose relevance depends on the physics of system and is often not clear a priori.
@@
As a representative case study, we focus on a prototypical molecular system with non-trivial internal dynamics. Specifically, we analyze atomistic Molecular Dynamics (MD) trajectories of a system consisting of 2048 TIP4P/ICE [abascal2005potential] molecules at the melting temperature [caruso_timesoap_2023]. Initially, half of the molecules are in the liquid phase, while the other half form solid ice. After equilibration, a $50$ ns production run was performed under $NPT$ conditions (see Methods section for complete details on the molecular model and simulation setup).
@@
Recent studies have demonstrated that a general yet relatively simple descriptor, the Local Environments and Neighbors Shuffling (LENS), can effectively capture the intrinsic dynamics of such system with remarkable sensitivity. As illustrated in Fig. 1A, the LENS value for every particle in the system is computed for each sampling interval (of length $\delta t$) along the MD trajectory. This is done by measuring the variation, frame by frame, in the string of the IDs of the neighbors within a specified cutoff radius (dashed circles in Fig. 1A). LENS captures events such as the addition, removal, or permutation of neighbors during $\delta t$, offering insights into the local dynamics of the environment surrounding each particle in the system.
@@
In this case, 2048 LENS time-series are extracted – one for each molecule, considering only the oxygen atoms – each consisting of 500 frames ($\delta t=100$ ps); see Fig. 1B. The probability kernel density estimate (KDE) of the data points reveals two distinct peaks (Fig. 1B, shown in blue). The first and most prominent peak occurs at LENS intensity $\sim 0.05$, corresponding to molecules that are nearly static. These represent the molecules in the ice phase, vibrating around their crystalline positions. The second, lower-density peak appears at a LENS intensity of $\sim 0.35$, corresponding to the more dynamic molecules in the liquid phase.
@@
These are the two main clusters that standard pattern recognition approaches, which analyze the entire dataset while ignoring time correlations, readily identify from the trajectories of this system [crippa_machine_2023, caruso_timesoap_2023]. However, recent studies have shown that additional, highly relevant information is hidden within these data, masked by the noise of these two dominant clusters [crippa_detecting_2023]. This information can be extracted by accounting for time correlations in the data of individual molecules, treating these data as time-series, and systematically distinguishing meaningful fluctuations from noise, using single-point time-series clustering algorithms, such as Onion Clustering [becchi_layer-by-layer_2024, martino2024data, lionello2024relevant].
@@
As shown in Fig. 1C, Onion Clustering performs a single-point time-series analysis on the trajectories of all molecules in the system, classifying them based on the similarity of their vibrational dynamics. This method employs an iterative detect-classify-archive approach with a resolution that adaptively increases at each iteration. Onion Clustering systematically repeats the clustering analysis across all possible temporal resolutions $\Delta t$, starting from the maximum resolution, which is twice the minimum time step (in this case, $200$ ps), and progressively reducing the resolution down to the minimum, corresponding to the total duration of the simulation ($50$ ns).
@@
The plot in Fig. 1D (blue curve) shows the number of statistically distinct environments identified as a function of the time resolution. The analysis reveals not just two but up to three distinct environments from the LENS time-series. Physically, these correspond to the ice (gray in Fig. 1E), liquid (blue), and the ice-water interface (red). From the Onion plot, it is evident that these three clusters are distinguishable at resolutions of $\Delta t\leq 10$ ns. Beyond this threshold, the resolution becomes insufficient to differentiate the ice-water interface as a dynamically distinct environment, resulting in a reduction from three clusters to two. This finding suggests that the characteristic residence time of water molecules in the interface is approximately 10 ns; using a lower resolution than this prevents the identification of the interface environment as distinct.
@@
Notably, for $\Delta t>20$ ns, only a single stable environment, corresponding to the ice bulk, can be identified. This indicates that some molecules within the ice bulk remain in this environment throughout the entire duration of the simulation. At these resolutions, $\Delta t>20$ ns, the analysis effectively detects one stable environment (ice), while all other data points are classified as part of an “unclassifiable data” cluster. These unclassified molecules, at this resolution, cannot be attributed to any stable environment, reflecting the inability of such a coarse temporal resolution to resolve their dynamic behavior. Notice that, as the value of $\Delta t$ increases, the fraction of unclassified data points increases as well (orange curve in Fig. 1D), until most of the data are in this category, because the time resolution is too coarse to resolve most of the relevant system dynamics.
@@
For more details regarding the descriptor and the clustering method, we refer the interested reader to Refs. [crippa_detecting_2023] and [becchi_layer-by-layer_2024]. A crucial strength of Onion Clustering lies in its ability to transform a potential limitation of clustering analysis – the dependence on the choice of time resolution – into an advantage. By eliminating the need for the user to select a specific time resolution a priori, Onion Clustering identifies the optimal resolution windows for studying the system’s physics. In this case, it reveals that resolutions of $\Delta t<10$ ns are ideal, as they allow the detection of three distinct clusters while minimizing the fraction of unclassifiable data points.
@@
It is worth highlighting that by repeating the analysis across various temporal resolutions and visualizing the outcomes, the approach becomes independent of any preconceived assumptions concerning the optimal temporal resolution for studying the system’s physics. Nonetheless, one key parameter still influences the analysis: the cutoff used in calculating the descriptor (LENS in this case, though this applies to many other descriptors as well). It is well known that varying the cutoff when using a descriptor can significantly impact the results. This raises an important question: what is the optimal spatial resolution (i.e., the best cutoff radius) for maximizing the extraction and classification of information from the data and ultimately the understanding of the system’s physics? While the answer to this fundamental question remains often elusive, the methodology presented here provides a robust framework to identify the “optimal spatiotemporal resolution” for analyzing a system solely based on the data extracted from its trajectories.
@@
Figure 2: Effect of spatial and temporal resolutions on information extraction from LENS time-series data. A Radial distribution function $g(r)$ of the oxygen atoms of all water molecules in the system: relevant $g(r)$ minima are detected (red circles) and used as critical cutoff radii $r_{c}$ to calculate LENS signals retaining information on local vs. non-local events/phenomena. B-G Onion plots obtained from the analyses of the LENS time-series with different $r_{c}$. The number of resolved micro-clusters is shown in blue, while the fraction of unclassifiable data (in the ENV0 cluster) is shown in orange. The red vertical dashed line shows the example time resolution of $\Delta t=1.1$ ns used for the snapshots in panel I. H In blue: mean number of classifiable micro-clusters (ENVs) before the fraction of unclassifiable data reaches 50%, as a function of the $r_{c}$ used in the analysis. The data show a clear trend (blue curve), where the maximum efficiency in information extraction-and-discretization is obtained at $r_{c}\sim 10$: namely, when accounting for events involving up to the 3rd-4th neighbors shell. I Representative MD snapshot of the system where the water molecules are colored according to the clustering obtained with the different $r_{c}$ in the analyses of panels B-G at the example time-resolution of $\Delta t=1.1$ ns: the red ice/water interface can be resolved in the spatial resolution range of $8.5\leq r_{c}\leq 15$, and for temporal resolutions higher than $\Delta t<10$ ns.

### v4

# Data-driven assessment of optimal spatiotemporal resolutions for information extraction in noisy time series data

@@ Abstract
In general, comprehension of any type of complex system depends on the resolution used to examine the phenomena occurring within it. However, identifying a priori, for example, the best time frequencies/scales to study a certain system over-time, or the spatial distances at which correlations, symmetries, and fluctuations are, most often non-trivial. Here we describe an unsupervised approach that, starting solely from the data of a system, allows learning the characteristic length scales of the dominant key events/processes and the optimal spatiotemporal resolutions to characterize them. We tested this approach on time series data obtained from simulation or experimental trajectories of various example many-body complex systems ranging from the atomic to the macroscopic scale and having diverse internal dynamic complexities. Our method automatically analyzes the system data by analyzing correlations at all relevant inter-particle distances and at all possible inter-frame intervals in which their time series can be subdivided, namely, at all space and time resolutions. The optimal spatiotemporal resolution for studying a certain system thus maximizes information extraction and classification from the system’s data, which we prove to be related to the characteristic spatiotemporal length scales of the local/collective physical events dominating it. This approach is broadly applicable and can be used to optimize the study of different types of data (static distributions, time series, or signals). The concept of “optimal resolution” has a general character and provides a robust basis for characterizing any type of system based on its data, as well as to guide data analysis in general.
@@ Introduction
Understanding the physics of complex many-body systems with rich internal dynamics is typically a challenging task ten Wolde and Frenkel (1997); Baletto (2019); Gasparotto et al. (2020); Bochicchio et al. (2019); de Marco et al. (2021); Cioni et al. (2023); Crippa et al. (2023a). This holds true across various scales Cho et al. (2021), from complex phenomena at the molecular level, such as nucleation events, phase transitions and ensemble dynamic adaptations in complex molecular systems ten Wolde and Frenkel (1997), to non-trivial collective behaviors in complex scale systems such as fish schools Porfiri (2018); Butail, Mwaffo, and Porfiri (2016), bird flocks Cavagna et al. (2010); Attanasi et al. (2014); Nagy et al. (2010) and socioeconomic or stock market systems Liu, Ungar, and Kording (2021); Mantegna and Stanley (1999); Duan et al. (2022). Typically, the first challenge in studying such complex systems is to extract relevant information about the natural dynamical behavior of their constituent units, a fundamental step in reconstructing the complex internal physics that characterize them Martino et al. (2024).
@@
The study of a complex system typically begins with the acquisition of trajectories, such as the positions and velocities of the individual units composing the system, obtained either through experiments or simulations. An efficient method is required to extract physically relevant information from these trajectories. A common approach involves using descriptors that convert raw trajectories into time series that can represent functions of positions (e.g. mutual arrangements) and / or velocities of the constituent units over time Kathirgamanathan and Cunningham (2020); Schmidt et al. (2023). The choice of descriptor is critical, as different descriptors enable the extraction of different types of physical information about the system’s behavior Caruso et al. (2023); Crippa et al. (2022); Martino et al. (2024). Among the various available descriptors, a distinction can be made between physics-inspired descriptors and "abstract" or general ones. Physics-inspired descriptors are tailored to study specific systems by leveraging the prior knowledge of their physics. Although they can be highly effective for investigating particular types of systems, they are typically less transferable to others. Furthermore, these descriptors may suffer from inherent biases and potentially overlook important and unforeseen information. To address these limitations, more general “abstract" descriptors – such as correlation functions or relative distances, to name a few commonly used examples – can be employed.
@@
More advanced and complex abstract descriptors, such as the Smooth Overlap of Atomic Positions (SOAP) Bartók, Kondor, and Csányi (2013) and Atomic Cluster Expansions (ACE) Drautz (2019) have also proven to be highly useful. Unlike system-specific descriptors, these general approaches are not tailored ad hoc to any particular system, making them applicable to virtually any system type. Their unbiased nature, stemming from the lack of reliance on prior knowledge of the system’s physics, allows them to be broadly applicable and transferable for studying a wide range of complex molecular systems. Pattern recognition approaches for analyzing datasets, such as those from ACE or SOAP, have successfully provided valuable insights into the physics of systems across scales, from molecular to macroscopic, including aqueous, metallic, and other complex systems Fitzner et al. (2019); Capelli, Muniz-Miranda, and Pavan (2022); Lionello et al. (2023); Cioni et al. (2023); Donkor, Laio, and Hassanali (2023); Cioni et al. (2023); Crippa et al. (2023b); Cioni et al. (2024); Donkor et al. (2024); Martino et al. (2024); Perrone et al. (2024); Cheng et al. (2019); Rapetti et al. (2023), and large-scale phenomena such as wave propagation Becchi, Fantolino, and Pavan (2024); Caruso et al. (2024). Data-mining approaches often rely on identifying high-density peaks (dominant microscopic structural motifs) in datasets, treating the data as an ensemble in which time correlations are ignored. This can potentially lead to loss of information Martino et al. (2024); Lionello et al. (2024).
@@
Conversely, it has been demonstrated that studying data, acquired over time series (thus accounting for their time-correlation) allows decouple temporal from spatial information, which may provide more rich information than typical pattern-recognition approaches, including details on rare events, local or collective events that originate and amplify dynamically over time, etc Caruso et al. (2023); Crippa et al. (2023a); Butler, Fijan, and Glotzer (2024); Crippa et al. (2023b); Becchi, Fantolino, and Pavan (2024); Cioni et al. (2024); Perrone et al. (2024); Martino et al. (2024); Caruso et al. (2024); Lionello et al. (2024); Husic and Pande (2018). For example, descriptors such as timeSOAP Caruso et al. (2023) or Local Environments and Neighbors Shuffling (LENS) Crippa et al. (2023a) track how much and in what way the environment surrounding each unit in the system changes over time. These approaches shift the focus to identifying relevant fluctuations in noisy time series. An efficient method for this is Onion Clustering, an unsupervised and parameter-free algorithm that systematically detects statistically significant fluctuations in noisy time series of any type. It classifies these fluctuations based on their similarity, determining the number of distinct microscopic domains (clusters) can be identified in a time series as a function of the length of the signal intervals used in the analysis. Recent results from Onion clustering have revealed non-trivial dependencies on the temporal resolution used in the analysis. These findings show that, for a given system studied with a specific descriptor, there exists an optimal temporal resolution that maximizes the information extracted from its time series data Becchi, Fantolino, and Pavan (2024); Martino et al. (2024); Lionello et al. (2024). The considerations for temporal resolution also apply to spatial resolution. More broadly, this raises an important question: what is the optimal resolution – both spatial and temporal – for capturing the physical behavior of a system? For instance, it is known that adjusting the cutoff for radial-based local descriptors (e.g., capturing different ranges of short- or long-range effects) can significantly impact analysis results. Similar to the findings from Onion Clustering regarding temporal resolution, is there an optimal spatial resolution for extracting meaningful physical information from trajectory data? Moreover, is there a reliable method to determine this optimal resolution for a given system?
@@
In this work, we address this fundamental problem by studying various prototypical systems, ranging from atomic and molecular scales to microscopic and macroscopic scales, which exhibit different complex internal dynamics. We systematically analyzed the information contained in the time series data of their constituent units as a function of the spatial and temporal resolution used in the analyses. We demonstrated that different types of systems have distinct optimal spatiotemporal resolutions that maximize the extraction of relevant information from their trajectories, depending on the local or collective nature of the physical events/processes that occur within them. The characteristic spatiotemporal length scales of the dominant physical events dominating the system are in this way revealed in a purely data-driven way, and without prior assumptions on the best resolutions to capture and characterize them.
@@
While identifying the optimal time and space resolutions to study a system in which little is known a priori is challenging, this work demonstrates how such information is intrinsically present in the system’s data and it provides an efficient way to extract it. From a fundamental point of view, we provide clear demonstrations of the impact of such a data driven approach in revealing the physics of complex systems (or phenomena) without relying on prior knowledge or preconceived assumptions. At the same time, because every set of data has its own optimal spatiotemporal resolutions that allow for maximizing information extraction, we obtain a general concept and an useful approach to guide and assist the analysis of virtually any type of data.
@@ Results
Figure 1: Extracting information from, e.g., LENS time series of ice/water dynamic coexistence simulation trajectories. A Scheme of example local dynamical events captured by the LENS descriptor (permutation, addition, or subtraction of neighbors). B LENS signals for all the water molecules (their oxygen atoms) in the system as a function of simulation time, and their cumulative distribution (KDE: on the right, in blue). C Same LENS time series data with the background colored based on the three main micro-clusters detected by Onion Clustering (using a time-resolution of $\Delta t=1.1$ ns). The gaussian LENS environments are shown in the KDE (right) as solid gaussian curves, the inter-cluster thresholds are indicated as dotted horizontal lines. D Output Onion plot. The curve in blue (primary $y$-axis) shows the number of clusters classifiable by Onion Clustering as a function of the time resolution $\Delta t$. The curve in orange (secondary $y$-axis) shows the fraction of unclassifiable data points (stored into a cluster named ENV0) as a function of $\Delta t$ (i.e., fraction of dynamical events occurring faster than the resolution of the analysis). The vertical red dashed line indicates the time-resolution $\Delta t=1.1$ ns used for the other panels (clusters and thresholds in panel C). E MD snapshot of the ice/water coexistence simulation showing the TIP4P/ICE molecules (left) and coloring them based on the clusters classified at the example time-resolution, which correspond to the bulk of ice (in gray), to the bulk of the liquid phase (in blue), and to the ice/water interface (in red). In white are the unclassifiable points: these are domains in the liquid which freeze and re-melt faster than $\Delta t=1.1$ ns Becchi, Fantolino, and Pavan (2024); Capelli, Muniz-Miranda, and Pavan (2022); Crippa et al. (2023a); Caruso et al. (2023). F Schematic representation of the solvation shells accounting for the first, second, third, etc., neighbor particles around a unit i in a generic system: using different cutoff radii means capturing different types of events (e.g., local vs. non-local) whose relevance depends on the physics of system and is often not clear a priori.
@@
As a representative case study, we focused on a prototypical molecular system with non-trivial internal dynamics. Specifically, we analyzed the atomistic Molecular Dynamics (MD) trajectories of a system consisting of 2048 TIP4P/ICE Abascal et al. (2005) molecules at the melting temperature Caruso et al. (2023). Initially, half of the molecules are in the liquid phase, while the other half form solid ice. After equilibration, a $50$ ns production run was performed under $NPT$ conditions (see Methods section for complete details on the molecular model and simulation setup).
@@
Recent studies have demonstrated that a general yet relatively simple descriptor, LENS, can effectively capture the intrinsic dynamics of such system with remarkable sensitivity. As illustrated in Fig. 1A, the LENS value for every particle in the system was computed for each sampling interval (of length $\delta t$) along the MD trajectory. This is done by measuring the variation, frame by frame, in the string of the IDs of the neighbors within a specified cutoff radius (dashed circles in Fig. 1A). LENS captures events such as the addition, removal, or permutation of neighbors during $\delta t$, offering insights into the local dynamics of the environment surrounding each particle in the system.
@@
In this case, 2048 LENS time series were extracted – one for each molecule, considering only the oxygen atoms – each consisting of 500 frames ($\delta t=100$ ps); see Fig. 1B. The probability kernel density estimate (KDE) of the data points revealed two distinct peaks (Fig. 1B, shown in blue). The first and most prominent peak occurs at LENS intensity $\sim 0.05$, corresponding to molecules that are nearly static. These represent the molecules in the ice phase, vibrating around their crystalline positions. The second, lower-density peak appears at a LENS intensity of $\sim 0.35$, corresponding to the more dynamic molecules in the liquid phase.
@@
These are the two main clusters that standard pattern recognition approaches, which analyze the entire dataset while ignoring time correlations, are readily identifed from the trajectories of this system Crippa et al. (2023b); Caruso et al. (2023). However, recent studies have shown that additional, highly relevant information is hidden within these data, masked by the noise of these two dominant clusters Crippa et al. (2023a). This information can be extracted by accounting for time correlations in the data of individual molecules, treating these data as time series, and systematically distinguishing meaningful fluctuations from noise, using single-point time series clustering algorithms, such as Onion Clustering Becchi, Fantolino, and Pavan (2024); Martino et al. (2024); Lionello et al. (2024).
@@
As shown in Fig. 1C, Onion Clustering performs a single-point time series analysis on the trajectories of all molecules in the system, classifying them based on the similarity of their vibrational dynamics. This method employs an iterative detect-classify-archive approach with a resolution that adaptively increases at each iteration. Onion Clustering systematically repeats the clustering analysis across all possible temporal resolutions $\Delta t$, starting from the maximum resolution, which is twice the minimum time step (in this case, $200$ ps), and progressively reducing the resolution down to the minimum, corresponding to the total duration of the simulation ($50$ ns).
@@
The plot in Fig. 1D (blue curve) shows the number of statistically distinct environments identified as a function of the time resolution. The analysis reveals not just two but up to three distinct environments from the LENS time series. Physically, these correspond to the ice (gray in Fig. 1E), liquid (blue), and the ice-water interface (red). From the Onion plot, it is evident that these three clusters are distinguishable at resolutions of $\Delta t\leq 10$ ns. Beyond this threshold, the resolution becomes insufficient to differentiate the ice-water interface as a dynamically distinct environment, resulting in a reduction from three clusters to two. This finding suggests that the characteristic residence time of water molecules in the interface is approximately 10 ns; using a lower resolution than this prevents the identification of the interface environment as distinct.
@@
Notably, for $\Delta t>20$ ns, only a single stable environment, corresponding to the ice bulk, can be identified. This indicated that some molecules within the ice bulk remain in this environment throughout the entire duration of the simulation. At these resolutions, $\Delta t>20$ ns, the analysis effectively detects one stable environment (ice), while all other data points are classified as part of an “unclassifiable data" cluster. These unclassified molecules, at this resolution, cannot be attributed to any stable environment, reflecting the inability of such a coarse temporal resolution to resolve their dynamic behavior. Notice that, as the value of $\Delta t$ increases, the fraction of unclassified data points increases as well (orange curve in Fig. 1D), until most of the data are in this category, because the time resolution is too coarse to resolve most of the relevant system dynamics.
@@
For more details regarding the descriptor and the clustering method, we refer the interested reader to Refs. Crippa et al. (2023a) and Becchi, Fantolino, and Pavan (2024). A crucial strength of Onion Clustering lies in its ability to transform a potential limitation of clustering analysis – the dependence on the choice of time resolution – into an advantage. By eliminating the need for the user to select a specific time resolution a priori, Onion Clustering identifies the optimal resolution windows for studying the system’s physics. In this case, the resolution of $\Delta t<10$ ns is ideal, because it allows the detection of three distinct clusters while minimizing the fraction of unclassifiable data points.
@@
It is worth highlighting that by repeating the analysis across various temporal resolutions and visualizing the outcomes, the approach becomes independent of any preconceived assumptions concerning the optimal temporal resolution for studying the system’s physics. Nonetheless, one key parameter still influences the analysis: the cutoff used in calculating the descriptor (LENS in this case, though this applies to many other descriptors as well). It is well known that varying the cutoff when using a descriptor can significantly impact the results. This raises an important question: what is the optimal spatial resolution (i.e., the best cutoff radius) for maximizing the extraction and classification of information from the data and ultimately the understanding of the system’s physics? While the answer to this fundamental question remains often elusive, the methodology presented here provides a robust framework to identify the “optimal spatiotemporal resolution" for analyzing a system solely based on the data extracted from its trajectories.
@@
Figure 2: Effect of spatial and temporal resolutions on information extraction from LENS time series data. A Radial distribution function $g(r)$ of the oxygen atoms of all water molecules in the system: relevant $g(r)$ minima are detected (red circles) and used as critical cutoff radii $r_{c}$ to calculate LENS signals retaining information on local vs. non-local events/phenomena. B-G Onion plots obtained from the analyses of the LENS time series with different $r_{c}$. The number of resolved micro-clusters is shown in blue, while the fraction of unclassifiable data (in the ENV0 cluster) is shown in orange. The red vertical dashed line shows the example time resolution of $\Delta t=1.1$ ns used for the snapshots in panel I. H In blue: mean number of classifiable micro-clusters (ENVs) before the fraction of unclassifiable data reaches 50%, as a function of the $r_{c}$ used in the analysis. The data show a clear trend (blue curve), where the maximum efficiency in information extraction-and-discretization is obtained at $r_{c}\sim 10$: namely, when accounting for events involving up to the 3rd-4th neighbors shell. I Representative MD snapshot of the system where the water molecules are colored according to the clustering obtained with the different $r_{c}$ in the analyses of panels B-G at the example time-resolution of $\Delta t=1.1$ ns: the red ice/water interface can be resolved in the spatial resolution range of $8.5\leq r_{c}\leq 15$, and for temporal resolutions higher than $\Delta t<10$ ns.